# A Systematic Literature Review of Spatio-Temporal Graph Neural Network Models for Time Series Forecasting and Classification

### v1

@@ 5.7.5 Results
The most common error metrics used in the evaluation of the models in the ”Generic” group are, in order, the mean absolute error (MAE), correlation coefficient (CORR), root mean square error (RMSE), root relative squared error (RRSE), mean absolute percentage error (MAPE), and mean squared error (MSE). Differently from other metrics, where lower values indicate higher accuracy, the correlation coefficient (CORR) quantifies the strength of the the linear relationship between predicted and actual values, with higher values indicating a better performance. This Results subsection compares the results obtained for the Electricity consumption, Exchange rate, Solar energy and Traffic datasets. The results of the proposed models and benchmarks on METR-LA and PEMS datasets were already reported with a in Subsec. 5.5.5.
@@
The accuracy of the various models varies depending on the studied forecasting horizon but, overall, one can notice that the GNN models perform better than the others, with the static and dynamic graph learning network (SDGL) [art_78] being the most accurate.
@@
Also in this case, GNN models are the best overall. The discrete graph structure learning for time series model (GTS), proposed in 2021 by Shang et al. in Ref. [GTS] and used as a benchmark in Ref. [art_78], demonstrates itself to have a superior performance. It is therefore a suitable benchmark for comparison.
@@
The results indicate that the dynamic spatiotemporal interactive graph neural network (DSTIGNN) model [art_74] and the static and dynamic graph learning network (SDGL) model [art_78] are the most accurate on the solar energy dataset, for shorter and longer time horizons, respectively.
@@
The accuracy of the models varies by horizon and metric. However, the best one in terms of RRSE over all horizons is the temporal decomposition enhanced graph neural network for multivariate time series forecasting (TDG4MSF) [art_81].
@@ 5.8 Other topics
This final subsection includes 20 papers that fall outside the previously discussed groups. The case studies include cellular traffic prediction, recommendation systems for user preferences, and other forecasting or classification problems. Given the heterogeneity of the papers, only a general overview is provided here, and no comparisons of the approaches or discussions of benchmarks are presented.
@@ 5.8.1 Overview
Since GNNs have recently gained considerable popularity, many recent papers focus on applications that do not relate to the categories discussed so far. A frequently discussed topic in this ”Other topics” category is cellular traffic prediction ([art_138], [art_144], [art_148]), in terms of SMS messages, calls, internet connections, phone’s traffic statistics or locations, which are essential for the cellular network resource management system. Another topic is modulation classification ([art_12], [art_13], [4]), which is the process of determining the modulation used at the transmitter based on observations of the received signal [modulation_classification_definition]. A further area of investigation is the study of user preferences ([art_147], [art_151]) to enable personalized experiences and recommendations.
Other applications discussed in the papers of this section are encrypted traffic classification [art_118], urban spatio-temporal event prediction [art_136], prediction of the quality of operational processes [art_137], seismic intensity forecasting [art_139] (with source code available at https://github.com/StefanBloemheuvel/GCNTimeseriesRegression), prediction of spatio-temporal dynamics with complex structures [art_140], tailing dam monitoring [art_141], forecasting of educational video engagement [art_142], subsurface production forecasting [art_143], time series prediction for data centers maintenance [art_145], hydraulic runoff [art_146], water demand forecasting [art_149], and human activity recognition [art_159] (with source code available at https://github.com/riktimmondal/HAR-Sensor).
@@ 5.8.2 Datasets
Among the papers selected here, there are only three shared datasets, two for modulation classification and one for cellular traffic prediction. However, the datasets are handled in different ways, either in terms of pre-processing techniques, selected time windows, or train/test split percentages. For this reason, the results of these papers are not directly comparable among each other.
The two datasets for modulation classification used by Refs. [art_12], [art_13] and [4] are RML2016.10a (https://www.kaggle.com/datasets/raindrops12/rml201610a, source paper: http://dx.doi.org/10.1007/978-3-319-44188-7_16) and RML2016.10b (https://www.kaggle.com/datasets/marwanabudeeb/rml201610b, source paper: https://pubs.gnuradio.org/index.php/grcon/article/view/11). They include 220,000 and 1.2 million samples respectively, in many modulation types. Each modulation type has 20 levels of signal-to-noise ratios at 2 dB intervals from -20 dB to 18 dB.
As for the cellular traffic prediction dataset used in Refs. [art_138] and [art_144], it was proposed by Telecom Italia and MIT Media Lab during the ”Telecom Italia Big Data Challenge”, and it is available at https://dataverse.harvard.edu/dataverse/bigdatachallenge. The investigated geographical area is divided in a grid, and each cell records the number of SMS messages, calls, and wireless network traffic data in a 10 minutes interval.
@@ 6 Discussion
The aim of this SLR is to provide a comprehensive and detailed overview of the use of spatio-temporal GNN models for time series classification and forecasting in various fields, and to assess their performance in different application domains. Existing literature has been collected and synthesized in order to present an overview of datasets, models, and tables of results to support researchers in their future work. In the Introduction two sets of research questions were defined: a generic set and a specific set. Now, the following discussion will answer those research questions. Actually, most of the general research questions were already addressed in the overview presented in Sec. 3. As for the tools, the GNN community mostly uses Python and PyTorch, although some researchers use TensorFlow. Unfortunately, only few papers provide a link to the source code of the proposed model. Almost all papers are funded by public or private entities.
As for the specific questions, the answers are the following.
SQ1 (Applications). The three most investigated fields are, in order, ”Mobility”, ”Environment”, and ”Generic”. Regarding the first two groups, their datasets can be naturally translated into graphs, which explains why they have been studied more extensively over time. In contrast, the ”Generic” group has recently gained a significant interest for the applicability to broader contexts. As for the differences between the various fields, the main ones are related to the definition of the graph, which is not always explicit, and the benchmarks used, often related to the mindset of specific communities. It is not easy to compare results across different applications and to determine the most promising fields. The studies in the ”Generic” group appear promising across a wide range of domains.
@@
SQ3 (Taxonomy). As observed in Ref. [survey_spatiotemporal], there are two main approaches to the design of spatio-temporal GNNs: one that treats the spatial and temporal components in separate modules, and one that integrates and processes them together. The analysis of the collected papers reveals that the most common approach in literature is that of separate modules. Specifically, researchers often address the spatial and temporal aspects of the problem independently from each other, and focus on each module separately. The proposed taxonomy of GNN models refers only to the spatial component, and the review indicates that the convolutional and the attentional approaches are the prevalent ones. Approximately 62% of the models are purely convolutional, 25% are purely attentional, and 8% are hybrid convolutional-attentional. With regard to the temporal component, the recurrent structures of GRUs and the attention mechanism are widely used.
@@
SQ4 (Benchmark models). As for the benchmark models, there are many options available, and their choice depends on the specific application. In fields such as energy, Finance, health, and predictive monitoring, the focus tends to be on simpler statistical and classical machine learning benchmarks, with only a limited use of GNN benchmarks. In the ”Generic” group, many recent machine learning benchmark models based on the Transformer architecture have emerged. Notably, in ”Mobility”, ”Environment”, and ”Generic” groups (which are also the most investigated fields), there are many reference GNN benchmark models. The most prevalent benchmarks are the graph convolutional neural network with long-short term memory (GC-LSTM) [GC-LSTM], spatio-temporal graph convolutional network (STGCN) [STGCN], attention-based spatial-temporal graph convolutional network (ASTGCN) [ASTGCN], diffusion convolutional recurrent neural network (DCRNN) [DCRNN], spatio-temporal graph convolutional network (STGCN) [STGCN], and multivariate time series forecasting with graph neural network (MTGNN) [MTGNN].
@@
SQ5 (Benchmark datasets). The datasets mentioned in the selected papers are closely related to the specific case of study. Even though some benchmark datasets are listed in the ”Generic” group, there is no common standard dataset for the entire GNN research community. It would be beneficial to agreeing on, and start adopting some of the most commonly used datasets such as the traffic ones, with the goal of developing a shared benchmark dataset for the whole research community. Hopefully, this review could serve this purpose by providing comprehensive tables of the results of the models on different datasets to facilitate comparison between spatio-temporal GNN models.
@@
SQ6 (Modeling paradigms). As for the modeling paradigm, most of the selected papers work with a homogeneous graph, which models the relationships among multiple entities of the same nature. This is due to the fact that many basic GNN algorithms were originally developed for homogeneous graphs, and because it is easier to identify relationships between quantities of the same type. Also, in many cases, the focus is on multivariate series where the variables of interest and the target quantities are inherently of the same nature.
@@
SQ7 (Metrics). Regarding the error metrics, their choice is highly dependent on the specific case study. However, the most common ones are mean absolute error (MAE), mean squared error (MSE), and mean absolute percentage error (MAPE) for forecasting problems, and accuracy for classification problems.
@@ 7 Limits, challenges, and future research directions
Comparability. As highlighted in Ref. [cit_rev2], the evaluation of GNN models has improved thanks to the introduction of the Open Graph Benchmark (OGB) [NEURIPS2020_fb60d411], which provides a standardized evaluation framework and a variety of benchmark graph datasets. However, there is currently no standardized benchmark for spatio-temporal GNNs. As a result, each model is evaluated on its own selection of datasets, and this fragmentation makes it difficult to compare the results of different studies. To alleviate this problem, this review presented all the information gathered from the selected papers, including datasets, benchmarks, codes, and tables of results. It is intended to serve as a detailed overview and a foundation for further exploration, in the hope that researchers will begin to examine the data collected in the presented tables and, over time, identify relevant datasets and benchmark models.
@@
Reproducibility. The limited availability of links to repositories, source code, and datasets makes it difficult to evaluate the progress of knowledge in spatio-temporal GNNs. Moreover, in many papers the authors do not give detailed information about the model and the graph structure, like definition of the graph, number of nodes, calculation of edge weights. This issues complicate the verification of results and the assessment of the reproducibility of experiments, which are essential to drive further research in this field. Without access to these fundamental resources, it is extremely difficult to assess the accuracy and significance of the proposed models.
Explainability. A crucial aspect of GNN models is explainability, intended as the ability to interpret and understand their decision-making processes. Despite their attempt to explicitly model spatial relationships between series, GNNs are often considered as ”black box” models. None of the selected papers adequately tackle the concept of explainability. However, recent literature is beginning to investigate this important topic. For instance, few papers studied perturbation-based explanation methods, which however have proven to be somewhat ineffective as reported in Ref. [interpret_3]. Other approaches, such as GNNExplainer [interpret_2], aim to identify a small subset of node features that have a crucial role in GNN’s predictions. Recently, some researchers have started to explore graph counterfactuals as a means of generating explanations, as reviewed in Ref. [interpret_1].
@@
Poor information capacity. Another limitation of GNNs lies in the poor information behind the design of the graph structure and the model definition. A poorly constructed graph with unrepresentative or overly sparse connections, as well as a lack of physical constraints in the equations of the model, can significantly affect the performance. Although efforts have been made to address these issues, such as integrating physical constraints through differential equations, these approaches are not always effective. Therefore, it is essential to develop new techniques to overcome these limitations.
Heterogeneity. The majority of current GNN models are designed to deal with homogeneous graphs, where nodes and edges are all of the same type. As a result, it is challenging to use these GNNs with heterogeneous graphs, which have different types of nodes and edges or different inputs. Therefore, further research is necessary to develop GNN models that can effectively capture the interactions between various types of nodes and edges.
@@
Scalability. Spatio-temporal GNNs are widely used for modeling and analyzing large and complex time series-based network structures. However, GNN models often require a significant amount of memory to compute the adjacency matrix and node embeddings, especially in the case of dynamic graph structures. This scalability challenge results in high computational costs, and in many cases the necessity to use significant GPU resources. Therefore, the development of more scalable GNN models is crucial to facilitate time series analysis even in environments with limited computing power.

### v2

@@ 5.7.5 Results
The most common error metrics used in the evaluation of the models in the “Generic” group are, in order, the mean absolute error (MAE), correlation coefficient (CORR), root mean squared error (RMSE), root relative squared error (RRSE), mean absolute percentage error (MAPE), and mean squared error (MSE). Differently from other metrics, where lower values indicate higher accuracy, the correlation coefficient (CORR) quantifies the strength of the the linear relationship between predicted and actual values, with higher values indicating a better performance. This subsection about results compares the results obtained for the Electricity consumption, Exchange rate, Solar energy and Traffic datasets. The results of the proposed models and benchmarks on METR-LA and PEMS datasets were already reported with a in Subsec. 5.5.5.
@@
The accuracy of the various models varies depending on the studied forecasting horizon but, overall, one can notice that the GNN models perform better than the others, with the static and dynamic graph learning network (SDGL) [art_78] being the most accurate. The SDGL model, proposed by Li et al. in 2023, first employs a static and dynamic graph learning module that incorporates graph regularization to control the smoothness, connectivity, and sparsity of the learned graph. The graph input then passes through gated temporal convolutional networks and graph convolution modules, which are equipped with residual and skip connections that deliver the information to the output layer.
@@
Also in this case, GNN models are the best. The so-called discrete graph structure learning for time series model (GTS), proposed in 2021 by Shang et al. in Ref. [GTS] and used as a benchmark in Ref. [art_78], demonstrates itself to have a superior performance. It is therefore a suitable benchmark for comparison. The GTS model was introduced by Shang et al. in 2021 for traffic and sensor network forecasting. It includes a probabilistic graph learning module that optimizes the mean performance over the graph distribution, and a recurrent graph convolutional module to generate the forecasts.
@@
The results indicate that the dynamic spatiotemporal interactive graph neural network (DSTIGNN) model [art_74] and the static and dynamic graph learning network (SDGL) model [art_78] are the most accurate on the solar energy dataset, for shorter and longer time horizons, respectively. The DSTIGNN model, introduced by Gao et al. in 2021, employs a hierarchical structure consisting of a spatio-temporal interactive learning module, a dynamic graph inference module, multiple spatio-temporal layers, and an output module. The spatio-temporal interactive learning module splits the time series into two subsequences to capture both temporal and spatial correlations. Then, the graph inference module learns a dynamic adjacency matrix for the graph. Once the graph structure is determined, a stack of graph convolutional layers, enhanced with residual connections, along with a convolutional output module, generate the final forecast. The description of the SDGL instead has already been provided in this subsection, in the discussion of the results on the Electricity consumption dataset.
@@
The accuracy of the models varies by horizon and metric. However, the best one in terms of RRSE over all horizons is the temporal decomposition enhanced graph neural network for multivariate time series forecasting (TDG4MSF) [art_81], proposed by Miao et al. in 2021. The TDG4MSF model is composed of four components: temporal decomposition enhanced representation learning for extracting short and long term periodic patterns, graph structure learning, gated graph neural network depending on linear transformation for representation learning, and MLP-based output layer.
@@ 5.8 Other topics
This final subsection includes 29 papers that fall outside the previously discussed groups. The case studies include cellular traffic prediction, recommendation systems for user preferences, and other forecasting or classification problems. Given the heterogeneity of the papers, only a general overview is provided here, and no comparisons of the approaches or discussions of benchmarks are presented.
@@ 5.8.1 Overview
The interest in GNNs has grown so extensively that the selected themes are actually not sufficient to cover all relevant application domains. However, since only a few papers would belong to each potential additional theme, they are grouped under the label “Other Topics”. A frequently discussed topic in this “Other topics” category is cellular traffic prediction [art_138, art_144, art_148, art_468, art_457], in terms of SMS messages, calls, internet connections, phone’s traffic statistics or locations, which are essential for the cellular network resource management system. Another topic included is modulation classification [art_12, art_13, art_303, art_459, art_463], which is the process of determining the modulation used at the transmitter based on observations of the received signal [modulation_classification_definition]. Two further areas of investigation to be added are the study of user preferences [art_147, art_151] to enable personalized experiences and recommendations, and anomalous IoT intrusion detection [4, art_460].
Other applications discussed in the papers of this subsection are encrypted traffic classification [art_118], urban spatio-temporal event prediction [art_136], prediction of the quality of operational processes [art_137], seismic intensity forecasting [art_139] (with source code available at https://github.com/StefanBloemheuvel/GCNTimeseriesRegression), prediction of spatio-temporal dynamics with complex structures [art_140], tailing dam monitoring [art_141], forecasting of educational video engagement [art_142], subsurface production forecasting [art_143], time series prediction for data centers maintenance [art_145], hydraulic runoff [art_146], water demand forecasting [art_149], human activity recognition [art_159] (with source code available at https://github.com/riktimmondal/HAR-Sensor ), crime prediction [art_465] (with source code available at https://github.com/ZJUDataIntelligence/HDM-GNN), polypropylene production data [art_455], and tungsten flotation process [art_458] (with source code available at https://github.com/ywmcsu/SASGNN).
@@ 5.8.2 Datasets
Among the papers selected here, there are only four shared datasets, two for modulation classification, one for cellular traffic prediction, and one for botnet detection in IoT network traffic. However, the datasets are handled in different ways, either in terms of pre-processing techniques, selected time windows, or train/test split percentages. For this reason, the results of these papers are not directly comparable to each other.
The two datasets for modulation classification used in Refs. [art_12, art_13, art_303, art_463] are RML2016.10a (https://www.kaggle.com/datasets/raindrops12/rml201610a, source paper: http://dx.doi.org/10.1007/978-3-319-44188-7_16) and RML2016.10b (https://www.kaggle.com/datasets/marwanabudeeb/rml201610b, source paper: https://pubs.gnuradio.org/index.php/grcon/article/view/11). They include 220,000 and 1.2 million samples respectively, in many modulation types. Each modulation type has 20 levels of signal-to-noise ratios at 2 dB intervals from -20 dB to 18 dB.
As for the cellular traffic prediction dataset used in Refs. [art_138, art_144, art_468], it was proposed by Telecom Italia and MIT Media Lab during the “Telecom Italia Big Data Challenge”, and it is available at https://dataverse.harvard.edu/dataverse/bigdatachallenge. The investigated geographical area is divided in a grid, and each cell records the number of SMS messages, calls, and wireless network traffic data in a 10 minutes interval.
Finally, the Bot-IoT dataset used in Refs. [4, art_460] and available at https://www.unsw.adfa.edu.au/unsw-canberra-cyber/cybersecurity/ADFA-NB15-Datasets/bot_iot.php, was created by designing a realistic network environment in the Cyber Range Lab of UNSW Canberra, and incorporates a combination of normal and botnet traffic.
@@ 6 Discussion
As stated in the Introduction, the aim of this SLR is to provide a comprehensive and detailed overview of spatio-temporal GNN models for time series classification and forecasting in different fields, and to assess their performance. Existing literature was collected and synthesized to provide an overview of datasets, models, and tables of results that will hopefully assist researchers in their future work. In the Introduction two sets of research questions were stated: a generic set and a specific set. This section will collect the answers to those research questions.
Most of the general overview questions (GQs) were addressed in the overview presented in Sec. 3. As for the tools, the GNN community mostly uses Python and PyTorch, although some researchers use TensorFlow. Unfortunately only few papers provide a link to the source code of the proposed model. Almost all papers are funded by public or private entities.
As for the specific questions (SQs), the answers are the following.
SQ1 (Applications). The three most investigated fields are “Mobility”, “Environment”, and “Generic”. Regarding the first two thematic groups, their datasets can be naturally translated into graphs, which explains why they have been studied more extensively over time. In contrast, the “Generic” thematic group has only recently gained a significant interest. As for the differences among the thematic groups, the main ones are related to the definition of the graph, which is not always explicit, and the benchmarks used, often related to the mindset of specific communities. It is not easy to compare results across different applications and to determine the most promising fields. The studies in the “Generic” thematic group appear promising across a wide range of domains.
@@
SQ3 (Taxonomy). As observed in Ref. [survey_spatiotemporal], there are two main approaches for the design of spatio-temporal GNNs: one that treats the spatial and temporal components in separate modules, and another that integrates and processes them together. The analysis of the collected papers reveals that the most common approach in literature is that of separate modules. Specifically, researchers very often address the spatial and temporal aspects of the problem independently from each other, and focus on each module separately. Hence, the proposed taxonomy of GNN models takes its main perspective from the point of view of the spatial component. The review indicates that the convolutional and the attentional approaches are the prevalent ones. Approximately 62% of the models are purely convolutional, 21% are purely attentional, and 9% are hybrid convolutional-attentional. With regard to the temporal component, the recurrent structures of GRUs and the attention mechanism are widely used.
@@
SQ4 (Benchmark models). As for the benchmark models, there are many options available, and their choice depends on the specific application. In thematic groups such as “Energy”, “Finance”, “Health”, and “Predictive monitoring”, the focus tends to be on simpler statistical and non-GNN machine learning benchmarks, with only a limited use of GNN benchmarks. In the “Generic” group, many recent machine learning benchmark models based on the Transformer architecture have emerged. Notably, in “Mobility”, “Environment”, and “Generic” groups (which are also the most investigated fields), there are many reference GNN benchmark models. Among the most prevalent benchmarks there are: the graph convolutional neural network with long-short term memory (GC-LSTM) [GC-LSTM], spatio-temporal graph convolutional network (STGCN) [STGCN], attention-based spatial-temporal graph convolutional network (ASTGCN) [ASTGCN], diffusion convolutional recurrent neural network (DCRNN) [DCRNN], spatio-temporal graph convolutional network (STGCN) [STGCN], and multivariate time series forecasting with graph neural network (MTGNN) [MTGNN].
@@
SQ5 (Benchmark datasets). The datasets mentioned in the selected papers are closely related to the specific case of study. Even though some benchmark datasets are listed in the “Generic” thematic group, there is no common standard dataset for the entire spatio-temporal GNN research community. Hence, it would be beneficial for the community to agree on, and start adopting, some of the most commonly used datasets, with the goal of developing a shared benchmark dataset. A possible common benchmark could be the traffic one. Hopefully, this review could serve this purpose by providing comprehensive tables of the results of the models on different datasets to facilitate comparison between spatio-temporal GNN models.
@@
SQ6 (Modeling paradigms). As for the modeling paradigm used in each paper, most of the selected papers work with a homogeneous graph, which models the relationships among multiple entities of the same nature. This is due to the fact that many basic GNN algorithms were originally developed for homogeneous graphs, and because it is easier to identify relationships between quantities of the same type. Also, in many cases, the focus is on multivariate series where the variables of interest and the target quantities are inherently of the same nature.
@@
SQ7 (Metrics). Regarding the error metrics used in the papers, their choice is highly dependent on the specific case study. However, the most common ones are mean absolute error (MAE), mean squared error (MSE), and mean absolute percentage error (MAPE) for forecasting problems, and accuracy for classification problems.
A meta-analysis of papers from different areas indicates that some thematic groups have seen more intense research on GNN models than others. For instance, the papers in the “Mobility” and the “Generic” groups present more complex models than others, this being likely due to the more extensive research conducted in these fields, as also evidenced by the larger number of published papers. In contrast, the “Finance” and “Health” thematic groups have less papers, and often have only a limited number of benchmarks available for comparison. Another notable example is the “Energy” group, which has a relatively large number of papers, particularly in 2024. However, the energy community tends to align more closely with traditional econometrics or non-GNN machine learning models, and the proposed GNN models in this field are generally less complex than those found in other groups. This suggests that some fields are currently more prone to the development of innovative and complex models, whereas others require a greater focus on adapting existing models to real-world problems. Furthermore, papers in the “Energy” group often test their models on proprietary dataset, which further complicates the development of models across the entire group. Consequently, we expect some form of cross-pollination between the communities, which is likely to reduce these differences over time.
A final aspect to highlight is the importance of reproducibility and comparability of these studies, which is essential for the advancement of scientific knowledge. In some fields, a greater attention to reproducibility can be observed, also thanks to the availability of datasets, a larger number of benchmarks, and more detailed descriptions of the models. Tab. 40 shows the average number of datasets and benchmark models used, as well as the percentage of papers reporting the source code for each group.
Table 40: Average number of datasets and benchmark models used, and percentage of papers with the link to the source code for each group.
Group Datasets (avg.) Benchmarks (avg.) Code availability
Energy 1.6 5.8 3%
Environment 1.3 6.5 26%
Finance 1.8 5.6 12%
Generic 4.2 10.0 41%
Health 1.7 7.7 53%
Mobility 2.6 8.4 27%
Other 2.1 5.6 14%
Predictive monitoring 2.1 8.1 14%
@@ 7 Limits, challenges, and future research directions
Comparability. As highlighted in Ref. [cit_rev2], the evaluation of GNN models has improved thanks to the introduction of the Open Graph Benchmark (OGB) [NEURIPS2020_fb60d411], which provides a standardized evaluation framework and a variety of benchmark graph datasets. However, there is currently no standardized benchmark for spatio-temporal GNNs. As a result, each model is evaluated on its own selection of datasets, and this fragmentation makes it difficult to compare the results of different studies. To alleviate this problem, this review presented all the information gathered from the selected papers, including datasets, benchmarks, codes, and tables of results. All that is intended to serve as a detailed overview and a foundation for further exploration, in the hope that researchers will begin to examine the data collected in the presented tables and, over time, identify relevant datasets and benchmark models. Moreover, we encourage the authors to discuss the relationship between complexity and performance of the proposed models to facilitate a more comprehensive comparison.
@@
Reproducibility. The limited availability of links to repositories, source code, and datasets makes it difficult to evaluate the progress of knowledge in spatio-temporal GNNs. Moreover, in many papers the authors do not give detailed information about the model and the graph structure, like definition of the graph, number of nodes, calculation of edge weights. This issue complicates the verification of results and the assessment of the reproducibility of experiments, which are essential to drive further research in this field. Without access to this fundamental information, it is extremely difficult to assess the accuracy and significance of the proposed models.
@@
Poor information capacity. Another limitation of GNNs lies in the limited amount of information used in the design of the graph structure and the model definition. A poorly constructed graph with unrepresentative or overly sparse connections, as well as a lack of physical constraints in the equations of the model, can significantly affect the performance. Although efforts have been made to address these issues, such as integrating physical constraints through differential equations, these approaches are not always effective. Therefore, it is essential to develop new techniques to overcome these limitations.
@@
Scalability. Spatio-temporal GNNs are widely used for modeling and analyzing large and complex time series-based network structures. However, GNN models often scale badly, and require a significant amount of memory and computing resources to calculate the adjacency matrix and node embeddings, especially in the case of dynamic graph structures. This challenge can result in high computational costs, and in many cases in the necessity of using intensively GPU devices. Therefore, the development of more scalable GNN models is crucial to facilitate time series analysis even in environments with limited computing power.
Heterogeneity. The majority of current GNN models are designed to deal with homogeneous graphs, where nodes and edges are all of the same nature. As a result, it would be challenging to use these GNNs with heterogeneous graphs which have different types of nodes and edges or different inputs. Therefore, further research is necessary to develop GNN models that can effectively capture the interactions between various types of nodes and edges.
Explainability. A crucial aspect for GNN models is explainability, intended as the ability to interpret and understand the decision-making processes implemented in them. Despite their attempt to explicitly model spatial relationships between series, GNNs are often considered as “black box” models. Remarkably, none of the selected papers adequately tackle the concept of explainability. Some models are claimed to be explainable because of the integration of non-linear physical laws (e.g., Ref. [art_401]), while other papers claim explainability on the basis of the possibility to visualize a graph (e.g., Ref. [art_360]) or because they present a feature aggregation form that can be intuitively understood. In addition, other approaches focus on determining feature importance through tools such as Captum [Captum], which enables the interpretation and understanding of PyTorch model predictions by exploring the features that contribute to a specific outcome (e.g., Ref. [art_361]). This highlights a research gap in this area. However, recent literature is beginning to investigate this important topic. For instance, few papers studied perturbation-based explanation methods, which however have proven to be somewhat ineffective, as reported in Ref. [interpret_3]. Other approaches, such as GNNExplainer [interpret_2], aim to identify a small subset of node features that have a crucial role in GNN’s predictions, as in Ref. [art_349, art_357]. Recently, some researchers have started to explore graph counterfactuals as a means of generating explanations, as reviewed in Ref. [interpret_1].
Quasi-equilibrium learning. As pointed out in Subsec. 4.2, GNNs were originally introduced in Ref. [Scarselli2009] by promoting an extension of the backpropagation algorithm which exploits a fixed-point principle. The subsequent introduction of the notion of convolutional graph has fundamentally changed such computational scheme. It is worth mentioning that in contemporary GNNs there is no need to wait for the system to reach equilibrium. For example, the deep neural network scheme no longer requires weight-sharing. These improvements have led to successful applications in many real-world problems, where modern features contribute significantly to both accuracy and computational efficiency. However, when analyzing the processing of temporal information, GNNs are often implemented within a single graph domain where a temporal function is defined. This is a case where the original computational approach proposed in Ref. [Scarselli2009] seems to be more adequate, especially when considering that the rate of weight updates is generally much slower than the rate of external temporal signals. This suggests the necessity of revisiting the process of convergence to the fixed point, and of exploring computational schemes which take place at quasi-equilibrium points. We expect that this approach can significantly alleviate the computational burden that is typical of the convergence to fixed-points required for spatio-temporal graph collections.